OpenStudy (lemanual):

Help me please A. Given the equation 5x − 4 = −2(3x + 2), solve for the variable. Explain each step and justify your process. B. Megan solved a similar equation below. Is Megan's solution correct? Explain why or why not. 3(2x − 4) = 5x − 1 6x − 12 = 5x − 1 11x − 12 = −1 11x = 11 x = 11 (10 points)

OpenStudy (jjwalton04):

Hold on

OpenStudy (lemanual):

Ok

OpenStudy (jjwalton04):

for A Let's solve your equation step-by-step. 5x−4=−2(3x+2) Step 1: Simplify both sides of the equation. 5x−4=−2(3x+2) Simplify: (Hide steps) Let's simplify step-by-step. 5x−4=−2(3x+2) 5x+−4=(−2)(3x)+(−2)(2)(Distribute) 5x+−4=−6x+−4 5x−4=−6x−4 Step 2: Add 6x to both sides. 5x−4+6x=−6x−4+6x 11x−4=−4 Step 3: Add 4 to both sides. 11x−4+4=−4+4 11x=0 Step 4: Divide both sides by 11. 11x/11 = 0 11 x=0 Answer: x=0

OpenStudy (jjwalton04):

Shes right Let's solve your equation step-by-step. 3(2x−4)=5x−1 Step 1: Simplify both sides of the equation. 3(2x−4)=5x−1 Simplify: (Hide steps) Let's simplify step-by-step. 3(2x−4)=5x−1 (3)(2x)+(3)(−4)=5x+−1(Distribute) 6x+−12=5x+−1 6x−12=5x−1 Step 2: Subtract 5x from both sides. 6x−12−5x=5x−1−5x x−12=−1 Step 3: Add 12 to both sides. x−12+12=−1+12 x=11 Answer: x=11

OpenStudy (lemanual):

Thanks guys I wish I can Best response both of you

OpenStudy (jjwalton04):

But carly her answer came out correct

OpenStudy (lemanual):

Ok thanks again

OpenStudy (carlyleukhardt):

Ok....Sorry

OpenStudy (lemanual):

Sorry for what?

OpenStudy (carlyleukhardt):

For being wrong it looked like she was wrong to me

OpenStudy (jjwalton04):

i just explained it. lol but you tried. After all im not a robot, i might be wrong

OpenStudy (lemanual):

Oh so megan is right

OpenStudy (jjwalton04):

Hehe a doggy xD

OpenStudy (carlyleukhardt):

I don't think so but he explained it better

OpenStudy (lemanual):

Ok

OpenStudy (lemanual):

Yes, the way she got the answer was wrong, even though she got the right answer.